### v5

@@ 1 Introduction
Consider the classic signalling game: a sender is informed about a payoff-relevant state of the world drawn from a known distribution and takes one of several possible actions; an uninformed receiver observes the action but not the state, and makes a decision. In a landmark article, Crawford and Sobel, (1982) (henceforth CS) showed that, even if the payoffs of both agents are independent of the sender’s action, there are equilibria where the action transmits information about the state, as long as the conflict of interest about the ideal receiver’s decision is not too large. By interpreting the payoff-irrelevant actions of the sender as “cheap talk”, CS delivers a powerful formal theory of communication. Non-committal and purely symbolic behaviour can convey information and help coordinate subsequent interactions even if rational agents do not share identical goals.
@@
In this paper, we compute stationary points of memoryless independent reinforcement learning algorithms playing the CS’s game of information transmission. [^1] Computational techniques are necessary because finding limit points of independent learning algorithms interacting with each other is, to date, an intractable mathematical problem. Methods relying on approximation via systems of differential equations are not directly applicable to our case (see Börgers and Sarin, (1997) and Banchio and Mantegazza, (2022)). These algorithms work roughly as follows. For each of a finite set of states, the sender keeps track of a vector, which stores its current estimates of the value of taking each action in that state. The receiver, instead, holds a vector for each of the signals the sender may send. Any such vector contains the receiver’s estimate of the value of each action following a given signal. In each period, the algorithms select actions following a softmax policy. Most likely, they take the highest-reward action according to their estimates, but with some probability they experiment with other actions. Such probability decays over time, depending on a hyperparameter (i.e., the temperature-decay factor). After both agents have moved, the relevant estimates are updated to account for the payoffs received. Another hyperparameter (i.e., the learning rate) establishes how much the current experience is weighted vis-a-vis the past. [^2] The reinforcement learning literature has proposed numerous learning algorithms. We chose one of the simplest forms of reinforcement learning available for our environment (Sutton and Barto,, 2018).
@@
We find that a sender and a receiver training together converge to Nash equilibrium behaviour with substantial information transmission, which grows monotonically as the bias decreases. Except at levels of the bias that make perfect information transmission an equilibrium, where behaviour is more nuanced, the mutual information between the distribution of the state and that of the message chosen by the sender (i.e., the informativeness of the sender’s cheap talk) matches that of the maximally informative, Pareto optimal, equilibrium or of the second most informative one. As the bias grows, the sender’s strict preference for the Pareto optimal equilibrium attenuates and, when near indifference is reached, information transmission drops to the second best equilibrium, which will itself become the optimal one once the bias increases further.
@@
Our algorithms play a very large number of identical one-shot games together. However, at convergence, play is independent of past choices. By design, algorithms are learning to play a static game, not a dynamic one. Therefore, the reader may wonder why not endowing them with the ability to respond to past history, as this allows for a richer set of outcomes and, quite possibly, more communication in the long run. While this would be a valid modelling choice were we interested in the emergence of bilateral communication, it is not the right approach to evaluate the emergence of language in a large population, which is the focus of our analysis and, arguably, the appropriate interpretation of equilibrium in CS. In fact, building on a standard, if not entirely uncontroversial argument in the theory of learning in games, the results we obtain should be interpreted as arising from learning within a large, relative to the discount factor, population of possibly more sophisticated reinforcers interacting anonymously (see Fudenberg and Kreps, (1993) and Fudenberg and Levine, (1998)).
@@
Our main contribution is to the theory of strategic information transmission. Following pioneering work in psychology (e.g., Bush and Mosteller, (1955)) and in game theory (e.g., Erev and Roth, (1998)), reinforcement learning offers a model of human behaviour alternative to the traditional game-theoretic one. [^3] Erev and Roth, (1998) wrote: “well-developed, cognitively informed adaptive game theory will complement conventional game theory, both as a theoretical tool and as a tool of applied economics.” In this light, our results complement the existing knowledge in two ways. First, we show that information transmission in cheap talk games is a robust feature of play, emerging from a different modelling approach to strategic interaction. As was hoped for by Erev and Roth, (1998), simple reinforcement learning algorithms seem to fit the experimental data obtained from human subjects better than equilibrium does. In fact, slight under-communication relative to the most informative equilibrium is a common feature of experimental implementations of the CS game involving human subjects. (see Dickhaut et al., (1995) and, especially, Cai and Wang, (2006)). Second, we contribute to the large game-theoretic literature on equilibrium selection in games with information transmission. As it is well known, such games possess many qualitatively different equilibria. Our main result delivers a cautionary tale regarding the consensus reached in the uniform-quadratic environment around the selection of the most informative and Pareto optimal equilibrium. In particular, in addition to convergence to the first best, we also show convergence to the second most informative equilibrium at certain levels of the bias, where such an equilibrium is also close to optimal for the sender. [^4] For an important recent contribution to the equilibrium selection literature see Chen et al., (2008). Most closely related to our work is perhaps the evolutionary and learning approach to selection, which shows that the most informative equilibrium is the evolutionarily stable outcome of the CS game when it exists (see Blume et al., (1993)) and the limit point of the best-response dynamics (see Gordon et al., (2022)).
@@
While experimental evidence shows that communication in cheap talk games with partial conflict of interests is achieved by humans (e.g., see Blume et al., (2020) for a survey), to our knowledge an analogous conclusion has not yet been robustly established for artificially intelligent agents (AI agents). Most of the machine learning literature has focused on games with common interest, observing that AI agents learn to communicate successfully (e.g., see Lazaridou et al., (2016), Havrylov and Titov, (2017), Foerster et al., (2016)). Instead, mostly negative results have been obtained in games where there is scope for information exchange but agents have conflicting interests (e.g., see Cao et al., (2018)). An important exception is Noukhovitch et al., (2021). They consider a CS game played on a circle, for which equilibrium characterisation is not available. Employing AI agents controlled by neural networks they show that communication is achieved even when the bias of the sender is non-zero. We depart from Noukhovitch et al., (2021) by employing simple reinforcement learners and by looking at the original (discretised) CS game. Doing this allows us to compare simulation outcomes to the theoretical benchmark and establish that communication often takes place at the highest level predicted by theory even when a very simple model of learning is adopted.
@@ 2 RL agents playing the cheap talk game
There are two agents, a sender ($S$) and a receiver ($R$). At the outset, a state $\theta$ is drawn from a known uniform distribution with PDF $p$ and support over a finite set $\Theta$, which is composed by $n$ uniformly spaced points in the interval $[0,1]$. The sender privately observes the realised state $\theta$ and sends a message $m\in M$ to the receiver, with $|M|=|\Theta|$. Then, the receiver observes message $m$ and takes an action $a\in A$, with $A$ formed by $12n-$ 1 uniformly spaced points in $[0,1]$. The receiver wants the action to match $\theta$. Her payoff is $u_{R}(\theta,a)=-(a-\theta)^{2}$. Given bias $b\in[0,\infty)$, the sender wants the action of the receiver to match $\theta+b$. Thus, his payoff is ${u_{S}(\theta,a)=-(a-\theta-b)^{2}}$. The bias parameter measures their divergence of interest.
A (Bayes) Nash equilibrium is defined as a state-dependent choice of messages for the sender and a message-dependent choice of actions for the receiver, such that there is no profitable deviation by either agent for any observed state or message. A babbling equilibrium always exists, where the sender’s message is independent of the state and the receiver plays her ex-ante optimal action.
We let two independent reinforcement learning agents play, as sender and receiver, the discretised cheap talk game. To allow learning, the two agents play the game multiple times, up to a maximum of $T=10^{7}$ periods. Both are programmed to take an action conditional on a state, first the sender and then the receiver. In each period, a state for the sender is drawn from $\Theta$ according to $p$, independently of previous interactions. Then, the sender takes an action from $M$, which represents the state for the receiver. Finally, the receiver takes an action from $A$ and agents collect their rewards. Because the underlying learning model is the same for both agents (i.e., both take an action conditional on a state), we describe it for a generic agent, with states and actions taking appropriate meaning based on who is playing.
@@ 3 Main Results
For our baseline case, we consider the discretised game with $n=6$ states in $[0,1]$, with any two adjacent states separated by a $0.2$ increment. Hence, $\Theta=\{0.0,0.2,0.4,0.6,0.8,1.0\}$, $M=\Theta$ and $A=\{0.0,0.1,\ldots,0.9,1.0\}$.
@@
We also compute the mutual information between state and message, normalised by the entropy of the state. This is equivalent to the expected relative reduction in the entropy of the state from knowing the message. Formally,
@@
This metric takes value 1 if knowledge of the message implies knowledge of the state, as in a perfectly informative equilibrium. It takes value 0 when state and message are statistically independent, as in the babbling equilibrium. Notably, in contrast to payoff measures of informativeness, the normalised mutual information in the uniform prior case weighs equally information transmitted about different parts of the state space.
@@ 3.1 Convergence to Nash
Figure 1: Top: Maximum probability mass that the policy of the sender (receiver) places on suboptimal messages (actions) across all states (and messages). Bottom: Potential ex-ante gain by a unilateral deviation across all states (and messages). Averages over $1000$ simulations.
@@
Figure 2: Frequency of simulations in which both agents place at most $0.01$ probability mass on suboptimal actions across all states. The horizontal dashed lines correspond to frequencies of 00 and $1$. Each graph corresponds to a different $(\lambda,\alpha)$ combination of hyperparameters.
@@ 3.2 Equilibrium Selection
The CS game of information transmission has multiple Nash equilibria and this remains the case in its discretised version. While babbling is the unique equilibrium when the bias is above 0.3, other equilibria that induce different levels of information transmission exist with lower bias. The existing literature does not offer a complete characterisation of such equilibria in the discretised model. Frug, (2016) (Proposition 2) shows that the set of Pareto efficient equilibria is a singleton, payoff-wise. This equilibrium, which Frug, (2016) characterises, represents an important benchmark for us and is henceforth referred to as “optimal”.
@@
Figure 3: Heathmap of the modal policies of sender (top) and receiver (top) for different levels of bias over 1000 independent simulations. All vertical pairs of strategies correspond to exact equilibria. Randomisation over messages is with equal probability as indicated by the same colour tone. Messages to the right of the dashed line are off the equilibrium path. To find the modal policy we relabeled messages in each simulation assigning a natural number to each message such that messages with smaller numbers are associated with smaller states.
@@
In Figure 4, we compare the distribution (blue heatmap) of ex-ante payoffs arising from the simulations to the theoretical bounds provided by the babbling equilibrium (dotted line) and the optimal equilibrium (in red) for the different levels of bias in the discretised $[0,0.5]$ interval. In Figure 6, we show the distribution of the normalised mutual information between state and message from our simulations, for the same range of biases. Finally, in Figure 6 we identify all monotone partitional equilibria that exist in our discretised game at different levels of the bias, distinguishing them by their level of informativeness. Equilibria where the algorithms converge are highlighted in blue.
@@
Figure 5: Normalised mutual information between the distribution of messages induced by the sender’s policy and the distribution of states of the world. The distribution over 1000 simulations is shown in shades of blue. The value associated with the optimal equilibrium is in red and the one associated with the worst equilibrium is dotted gray.
@@
In addition to showing robustness of the results, Figure 7 highlights that letting agents explore more extensively yields outcomes that are progressively closer to the ex-ante optimal equilibrium. The same trend naturally extends to the normalised mutual information between messages and states of the world.
@@ 4 Comparative statics and Robustness
In this section, we demonstrate that communication emerges robustly in information transmission games played by AI agents, beyond the classic uniform-quadratic case studied so far. First, we look at making the language more or less expressive than it is necessary to achieve all equilibria of the game and, analogously, the set of actions more or less large. Second, we report the results of simulations obtained for a variety of alternative assumptions on the information transmission game. We consider a higher and lower number of states, non-uniform distributions of the state, and utility functions that are not linear-quadratic. Throughout, we keep fixed the reinforcement learning hyperparameters as in our baseline configuration. We show for each case how the ex-ante expected reward of the agents is distributed over 1000 simulations.
@@ 4.1 Comparative statics on messages and actions
In Figure 9 we endow the sender with more or fewer messages than the number required to achieve full information transmission. In particular, while keeping fixed the number of states to 6, we look at the case where only 3 messages are available and the case where 9 messages are present. Results are in line with expectations. On the one hand, increasing the number of messages does not have a tangible effect on the qualitative conclusions reached in the previous section. The sender simply learns to avoid redundant messages and the receiver responds to those in a way that does not stimulate the sender to play those actions. Learning becomes slower though, because the dimensions of the Q-matrices increase. On the other hand, making the language less expressive may impede players from playing the more informative equilibria. In fact, when the constraint on messages is non-binding, behaviour is as in the baseline scenario and less noisy because less learning is required. Instead, when the constraint is binding, sender and receiver tend to set on the most informative equilibrium compatible with the number of messages available.

### v6

@@ 1 Introduction
Consider the classic signalling game: a sender is informed about a payoff-relevant parameter drawn from a known distribution and takes one of several possible actions; an uninformed receiver observes the sender’s action but not the parameter, and makes a decision. In a landmark article, Crawford and Sobel, (1982) (henceforth CS) showed that, even if the payoffs of both agents are independent of the sender’s action, there are equilibria where the action transmits information about the parameter, as long as the conflict of interest about the ideal receiver’s decision is not too large. By interpreting the payoff-irrelevant actions of the sender as “cheap talk”, CS delivers a powerful formal theory of communication. Non-committal and purely symbolic behaviour can convey information and help coordinate subsequent interactions even if rational agents do not share identical goals.
@@
In this paper, we compute stationary points of memoryless independent reinforcement learning algorithms playing the CS’s game of information transmission. [^1] Computational techniques are necessary because finding limit points of independent learning algorithms interacting with each other is, to date, an intractable mathematical problem. Methods relying on approximation via systems of differential equations are not directly applicable to our case (see Börgers and Sarin, (1997) and Banchio and Mantegazza, (2022)). These algorithms work roughly as follows. For each of a finite set of types, the sender keeps track of a vector, which stores its current estimates of the value of taking each action given the type. The receiver, instead, holds a vector for each of the signals the sender may send. Any such vector contains the receiver’s estimate of the value of each action following a given signal. In each period, the algorithms select actions following a softmax policy. Most likely, they take the highest-reward action according to their estimates, but with some probability they experiment with other actions. Such probability decays over time, depending on a hyperparameter (i.e., the temperature-decay factor). After both agents have moved, the relevant estimates are updated to account for the payoffs received. Another hyperparameter (i.e., the learning rate) establishes how much the current experience is weighted vis-a-vis the past. [^2] The reinforcement learning literature has proposed numerous learning algorithms. We chose one of the simplest forms of reinforcement learning available for our environment (Sutton and Barto,, 2018).
@@
We find that a sender and a receiver playing together in repeated instances of the CS’s game converge to Nash equilibrium behaviour with substantial information transmission, which grows monotonically as the bias decreases. Except at levels of the bias that make perfect information transmission an equilibrium, where behaviour is more nuanced, the mutual information between the distribution of the type and that of the message chosen by the sender (i.e., the informativeness of the sender’s cheap talk) matches that of the maximally informative, Pareto optimal, equilibrium or of the second most informative one. As the bias grows, the sender’s strict preference for the Pareto optimal equilibrium attenuates and, when near indifference is reached, information transmission drops to the second-best equilibrium, which itself becomes the optimal one once the bias increases further.
@@
Our algorithms play a very large number of identical one-shot games together. However, at convergence, play is independent of past choices. By design, algorithms are learning to play a static game, not a dynamic one. Therefore, the reader may wonder why not endowing them with the ability to respond to past history, as this allows for a richer set of outcomes and, quite possibly, more communication in the long run. While this would be a valid modelling choice were we interested in the emergence of bilateral communication, it is not the right approach to evaluate the emergence of language in a large population, which is the focus of our analysis and, arguably, the appropriate interpretation of equilibrium in CS. In fact, building on a standard, if not entirely uncontroversial argument in the theory of learning in games, our results should be interpreted as arising from learning within a large population of possibly more sophisticated reinforcers interacting anonymously (see Fudenberg and Kreps, (1993) and Fudenberg and Levine, (1998)).
@@
Our main contribution is to the theory of strategic information transmission. Following pioneering work in psychology (e.g., Bush and Mosteller, (1955)) and in game theory (e.g., Erev and Roth, (1998)), reinforcement learning offers a model of human behaviour alternative to the traditional game-theoretic one. [^3] Erev and Roth, (1998) wrote: “well-developed, cognitively informed adaptive game theory will complement conventional game theory, both as a theoretical tool and as a tool of applied economics.” However, in contrast to what was hoped for by Erev and Roth, (1998), simple reinforcement learning algorithms do not fit the experimental data well in this case. In fact, over-communication relative to the most informative equilibrium is a common feature of experimental implementations of the CS game involving human subjects. (see Dickhaut et al., (1995) and, especially, Cai and Wang, (2006)). In this light, our results complement the existing knowledge in two ways. First, we show that information transmission in cheap talk games with conflict of interest is a robust feature of play, emerging from a different modelling approach to strategic interaction. To our knowledge, theoretical results on the convergence of reinforcement learning algorithms to informative equilibria are only available in the case where agents have aligned interests. Both Hu et al., (2011) and Wärneryd, (1993) demonstrate convergence of forms of reinforcement learning algorithms to the most informative equilibrium. Second, we contribute to the large game-theoretic literature on equilibrium selection in games with information transmission. As it is well known, such games possess many qualitatively different equilibria. Our main result delivers a cautionary tale regarding the consensus reached in the uniform-quadratic environment around the selection of the most informative and Pareto optimal equilibrium. In particular, in addition to convergence to the first best, we also show convergence to the second most informative equilibrium at certain levels of the bias, where such an equilibrium is also close to optimal for the sender. [^4] For an important recent contribution to the equilibrium selection literature see Chen et al., (2008). Most closely related to our work is perhaps the evolutionary and learning approach to selection, which shows that the most informative equilibrium is the evolutionarily stable outcome of the CS game when it exists (see Blume et al., (1993)) and the limit point of the best-response dynamics (see Gordon et al., (2022); Sémirat and Forges, (2024)).
@@
While experimental evidence shows that communication in cheap talk games with partial conflict of interests is achieved by humans (e.g., see Blume et al., (2020) for a survey), to our knowledge an analogous conclusion has not yet been robustly established for artificially intelligent agents (AI agents). Most of the machine learning literature has focused on games with common interest, observing that AI agents learn to communicate successfully (e.g., see Lazaridou et al., (2016), Havrylov and Titov, (2017), Foerster et al., (2016)). Instead, mostly negative results have been obtained in games where agents have conflicting interests (e.g., see Cao et al., (2018)). An important exception is Noukhovitch et al., (2021). They consider a CS game played on a circle, for which equilibrium characterisation is not available. Employing AI agents controlled by neural networks they show that communication is achieved even when the bias of the sender is non-zero. We depart from Noukhovitch et al., (2021) by employing simple reinforcement learners and by looking at the original (discretised) CS game. Doing this allows us to compare simulation outcomes to the theoretical benchmark and establish that communication often takes place at the highest level predicted by theory even when a very simple model of learning is adopted.
@@ 2 RL agents playing the cheap talk game
There are two agents, a sender ($S$) and a receiver ($R$). At the outset, a type $\theta$ is drawn from a known uniform distribution with PDF $p$ and support over a finite set $\Theta$, which is composed by $n$ uniformly spaced points in the interval $[0,1]$. The sender privately observes the realised type $\theta$ and sends a message $m\in M$ to the receiver, with $|M|=|\Theta|$. Then, the receiver observes message $m$ and takes an action $a\in A$, with $A$ formed by $12n-$ 1 uniformly spaced points in $[0,1]$. The receiver wants the action to match $\theta$. Her payoff is $u_{R}(\theta,a)=-(a-\theta)^{2}$. Given bias $b\in[0,\infty)$, the sender wants the action of the receiver to match $\theta+b$. Thus, his payoff is ${u_{S}(\theta,a)=-(a-\theta-b)^{2}}$. The bias parameter measures their divergence of interest.
Following CS, a (Bayes) Nash equilibrium is a family of type-conditional probability distributions over messages for the sender and a choice of action conditional on the message for the receiver, such that there is no profitable deviation by the receiver or by any type of sender. Babbling equilibria always exists, where the sender’s message conveys no relevant information on the type and the receiver plays her ex-ante optimal action for messages that are sent with positive probability.
We let two independent reinforcement learning agents play, as sender and receiver, the discretised cheap talk game. To allow learning, the two agents play the game multiple times, up to a maximum of $T=10^{7}$ periods. Both are programmed to take an action conditional on a state, first the sender and then the receiver. In each period, a state for the sender is a type drawn from $\Theta$ according to $p$, independently of previous interactions. Then, the sender takes an action from $M$, which represents the state for the receiver. Finally, the receiver takes an action from $A$ and agents collect their rewards. Because the underlying learning model is the same for both agents (i.e., both take an action conditional on a state), we describe it for a generic agent, with states and actions taking appropriate meaning based on who is playing.
@@ 3 Main Results
For our baseline case, we consider the discretised game with $n=6$ types in $[0,1]$, with any two adjacent types separated by a $0.2$ increment. Hence, $\Theta=\{0.0,0.2,0.4,0.6,0.8,1.0\}$, $M=\Theta$ and $A=\{0.0,0.1,\ldots,0.9,1.0\}$.
@@
We also compute the mutual information between type and message, normalised by the entropy of the type. This is equivalent to the expected relative reduction in the entropy of the type from knowing the message. Formally,
@@
This metric takes value 1 if knowledge of the message implies knowledge of the type, as in a perfectly informative equilibrium. It takes value 0 when type and message are statistically independent, as in the babbling equilibrium. Other measures of informativeness do not have qualitatively different behaviour. In particular, in the uniform-quadratic case, the negative of the residual variance of the sender’s type given the message coincides with the ex-ante payoff of the receiver, $U_{R}$.
@@ 3.1 Convergence to Nash
Figure 1: Top: Maximum probability mass that the policy of the sender (receiver) places on suboptimal messages (actions) across all types (and messages). Bottom: Potential ex-ante gain by a unilateral deviation across all types (and messages). Averages over $1000$ simulations.
@@
Figure 2: Frequency of simulations in which both agents place at most $0.01$ probability mass on suboptimal actions across all states, for different levels of bias in $[0,0.5]$. Each graph has bias on the horizontal axis and frequency on the vertical axis, and corresponds to a specific $(\lambda,\alpha)$ combination of hyperparameters. Horizontal dashed lines indicate frequencies of 0 and 1.
@@ 3.2 Equilibrium Selection
The CS game of information transmission has multiple Nash equilibria and this remains the case in its discretised version. While babbling is the unique equilibrium when the bias is above 0.3, other equilibria that induce different levels of information transmission exist with lower bias. The existing literature does not offer a complete characterisation of such equilibria in the discretised model. Frug, (2016) (Proposition 2) shows that in the uniform-quadratic case the set of Pareto efficient equilibria is a singleton, payoff-wise. This equilibrium, which Frug, (2016) characterises, represents an important benchmark for us and is henceforth referred to as “optimal”.
@@
Figure 3: Heathmap of the modal policies of sender (top) and receiver (top) for different levels of bias over 1000 independent simulations. All vertical pairs of strategies correspond to exact equilibria. Randomisation over messages is with equal probability as indicated by the same colour tone. Messages to the right of the dashed line are off the equilibrium path. To find the modal policy we relabeled messages in each simulation assigning a natural number to each message such that messages with smaller numbers are associated with smaller types.
@@
In Figure 4, we compare the distribution (blue heatmap) of ex-ante payoffs arising from the simulations to the theoretical bounds provided by the babbling equilibrium (dotted line) and the optimal equilibrium (in red) for the different levels of bias in the discretised $[0,0.5]$ interval. In Figure 6, we show the distribution of the normalised mutual information between type and message from our simulations, for the same range of biases. Finally, in Figure 6 we identify all monotone partitional equilibria that exist in our discretised game at different levels of the bias, distinguishing them by their level of informativeness. Equilibria where the algorithms converge are highlighted in blue.
@@
Figure 5: Normalised mutual information between the distribution of messages induced by the sender’s policy and the distribution of sender’s types. The distribution over 1000 simulations is shown in shades of blue. The value associated with the optimal equilibrium is in red and the one associated with the worst equilibrium is dotted gray.
@@
In addition to showing robustness of the results, Figure 7 highlights that letting agents explore more extensively yields outcomes that are progressively closer to the ex-ante optimal equilibrium. The same trend naturally extends to the normalised mutual information between messages and types.
@@ 4 Comparative statics and Robustness
In this section, we demonstrate that communication emerges robustly in information transmission games played by AI agents, beyond the classic uniform-quadratic case studied so far. First, we look at making the language more or less expressive than it is necessary to achieve all equilibria of the game and, analogously, the set of actions more or less large. Second, we report the results of simulations obtained for a variety of alternative assumptions on the information transmission game. We consider a higher and lower number of types, non-uniform prior, and utility functions that are not linear-quadratic. Throughout, we keep fixed the reinforcement learning hyperparameters as in our baseline configuration. We show for each case how the ex-ante expected reward of the agents is distributed over 1000 simulations.
@@ 4.1 Comparative statics on messages and actions
In Figure 9 we endow the sender with more or fewer messages than the number required to achieve full information transmission. In particular, while keeping fixed the number of types to 6, we look at the case where only 3 messages are available and the case where 9 messages are present. Results are in line with expectations. On the one hand, increasing the number of messages does not have a tangible effect on the qualitative conclusions reached in the previous section. The sender simply learns to avoid redundant messages and the receiver responds to those in a way that does not stimulate the sender to play those actions. Learning becomes slower though, because the dimensions of the Q-matrices increase. On the other hand, making the language less expressive may impede players from playing the more informative equilibria. In fact, when the constraint on messages is non-binding, behaviour is as in the baseline scenario and less noisy because less learning is required. Instead, when the constraint is binding, sender and receiver tend to set on the most informative equilibrium compatible with the number of messages available.